# The Quest for the Right Mediator: Surveying Mechanistic Interpretability Through the Lens of Causal Mediation Analysis

### v2

@@ 7.2.2 Inherently interpretable models
More ambitiously, one could consider building models with inherently interpretable components—i.e., whose fundamental units of computation (or some subset thereof) are designed to be sparse, monosemantic, and/or human-interpretable, but ideally still expressive enough to attain good performance on downstream tasks. Examples based in neural networks include differentiable masks (De Cao et al., 2020; Bastings, Aziz, and Titov, 2019), transcoders (Dunefsky, Chlenski, and Nanda, 2024), codebook features (Tamkin, Taufeeque, and Goodman, 2023), and softmax linear units (Elhage et al., 2022a). These are primarily post hoc methods that decompose model components into interpretable units, but they could potentially be integrated into the network itself during pre-training alongside a loss term (in addition to the language modeling loss) that enables fine-grained interpretability at all stages of pre-training.
Alternatively, more focus could be devoted to building models that are designed from the ground up to be interpretable, such as backpack language models (Hewitt et al., 2023), concept bottleneck models (Koh et al., 2020; Oikarinen et al., 2023), or decision trees (Hu, Rudin, and Seltzer, 2019). A related idea is to train the model using loss terms that encourage success on intermediate tasks, or induce particular kinds of feature representations (Hupkes, Zuidema et al., 2017). Perhaps least invasively, we could consider pre-training methods that softly encourage interpretable features to be aligned to neuron bases; this would remove the need for optimization to find non-basis-aligned components, and therefore make interpreting model decisions significantly easier and less confounded. However, this would reduce the number of features that could be encoded per neuron, so it would likely require significantly more parameters, or accepting degradations in performance. Regardless, we believe that this line of work will improve our ability to explain the behaviors of language models via directly improving the human-understandability of its intermediate computations; this will make it far easier to explain model behaviors, verify mechanistic hypotheses, and localize/edit particular computations—but at the potential expense of performance on downstream tasks.

### v3

@@ 7.2.2 Inherently interpretable models
More ambitiously, one could consider building models with inherently interpretable components—i.e., whose fundamental units of computation (or some subset thereof) are designed to be sparse, monosemantic, and/or human-interpretable, but ideally still expressive enough to attain good performance on downstream tasks. Examples based in neural networks include differentiable masks (De Cao et al., 2020; Bastings, Aziz, and Titov, 2019), transcoders (Dunefsky, Chlenski, and Nanda, 2024), codebook features (Tamkin, Taufeeque, and Goodman, 2023), and softmax linear units (Elhage et al., 2022a). These are primarily post hoc methods that decompose model components into interpretable units, but they could potentially be integrated into the network itself during pre-training alongside a loss term (in addition to a typical cross-entropy loss) that enables fine-grained interpretability at all stages of pre-training.
Alternatively, more focus could be devoted to building models that are designed from the ground up to be interpretable, such as backpack language models (Hewitt et al., 2023), concept bottleneck models (Koh et al., 2020; Oikarinen et al., 2023), or decision trees (Hu, Rudin, and Seltzer, 2019). A related idea is to train the model using loss terms that encourage success on intermediate tasks, or induce particular kinds of feature representations (Hupkes, Zuidema et al., 2017). Perhaps least invasively, we could consider pre-training methods that softly encourage interpretable features to be aligned to neuron bases; this would remove the need for optimization to find non-basis-aligned components, and therefore make interpreting model decisions significantly easier and less confounded. However, this would reduce the number of features that could be encoded per neuron, so it would likely require significantly more parameters, or accepting degradations in performance. Regardless, we believe that this line of work will improve our ability to explain the behaviors of contemporary NN-based systems like neural language models via directly improving the human-understandability of their intermediate computations; this will make it far easier to explain model behaviors, verify mechanistic hypotheses, and localize/edit particular computations—but at the potential expense of performance on downstream tasks.